OpenStudy (danieke):

Fan and Medal On the total amount of money spent by a school district for 1980 through 2000,the percent p spent on athletics can be modled by this equation p=2.94(1.07)t use the formula to calculate how much the district spent on athletics in 1999.Round your answer to the nearest thousandth

OpenStudy (danieke):

@automaticloveletter

OpenStudy (automaticloveletter):

What are your choices just so that I can make sure it's the right answer?

OpenStudy (danieke):

theres no choices it free response

OpenStudy (automaticloveletter):

Hm, well this site gives you step by step on how to find the answer, wanna see?

OpenStudy (automaticloveletter):

Slope = 0.636/2.000 = 0.318 p-intercept = 0/5000 = 0.00000 t-intercept = 0/-15729 = -0.00000 http://www.tiger-algebra.com/drill/p=2.94%281.07%29t/